OpenStudy (anonymous):

An inequality is shown below: -np - 5 ≤ 4(c - 2) Which of the following solves for n? n ≥ - the quantity 4 times c minus 3 all over p n ≥ - the quantity 4 times c minus 13 all over p n ≤ - the quantity 4 times c minus 3 all over p n ≤ - the quantity 4 times c minus 13 all over p

OpenStudy (anonymous):

@rs00752

OpenStudy (anonymous):

is it (minus)np-5?

OpenStudy (anonymous):

its negative np I think

OpenStudy (anonymous):

and it is also (minus)- the quantity....

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

then first option. -np-5<4c-8 =>-np<4c-3 =>n>-(4c-3)/p

OpenStudy (anonymous):

thanks can I ask you 3 more questions and then im done for now

OpenStudy (anonymous):

ofcourse..it will also help me strenghten my concepts..

OpenStudy (anonymous):

haha cool okay 1st one is A recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 5, 1, 7, 13, 19, ... 5, -1, -7, -13, -19, ... 5, 11, 17, 23, 29, ... 5, -11, -17, -23, -29, ... second one Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1 third one What is the equation of the line that passes through the points (-1, 3) and (1, 11)? 4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7 that's it

OpenStudy (anonymous):

First question f(2)=f(2-1)-6=5-6=-1 f(3)=f(3-1)-6=-1-6=-7 f(4)=f(4-1)-6=-7-6=-13 so, the answer is 5,-1,-7,-13,-19...

OpenStudy (anonymous):

okay next hurry I have to leave soon lol

OpenStudy (anonymous):

third answer is last option

OpenStudy (anonymous):

second answer is 3rd opt

OpenStudy (anonymous):

i hav some doubt regarding secon question's answer

OpenStudy (anonymous):

wait so third answer is for number 3 and the second answer is for number 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

third answer for third question

OpenStudy (anonymous):

and as always please let me know if i did something wrong..

OpenStudy (anonymous):

the answer to second question is right